x+y-2z=5
x +2=4
y -z=6

Answer by t0hierry(194)

Equation 2 gives you x = 2
Equation 1 gives you y-2z = 3
Equation 3 gives you y-z = 6
subtract
-z = 3-6
or z = 3
y = 9
check back
x + y - 2z = 2 + 9 - 6 = 5
y - z = 9 -3 = 6